# Boundary of the Moduli Space of Stable Cubic Fivefolds

### v6

@@ 5.14 Case $k=14$
Let $X_{14}:=V(\phi^{\mathrm{nf}}_{14})\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
@@
where $L_{01}=V(x_{2},x_{3},x_{4},x_{5},x_{6})$ and $C$ is a $(2,2)$-complete intersection in $V(x_{0},x_{1},x_{2})\simeq\mathbb{P}^{3}$. We have $L_{01}\cap C=\varnothing$.
@@ 5.15 Case $k=15$
Let $X_{15}:=V(\phi^{\mathrm{nf}}_{15})\subset\mathbb{P}^{6}$. Then the singular locus consists of three irreducible components:
@@
where $C_{1}=V(x_{1},x_{3},x_{6},2x_{4}^{2}+x_{0}x_{5},3x_{2}^{2}+x_{0}x_{4})$, $C_{2}=V(x_{2},x_{4},x_{5},x_{3}^{2}+x_{0}x_{6},x_{1}^{2}+x_{0}x_{3})$, and $L=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C_{1}$ and $C_{2}$ are curves of degree $4$ (complete intersections of two quadrics in $\mathbb{P}^{3}$), and $L$ is a line.
@@ 5.17 Case $k=17$
Let $X_{17}:=V(\phi^{\mathrm{nf}}_{17})\subset\mathbb{P}^{6}$. Then the singular locus consists of two disjoint components:
@@
where $C=V(x_{3},x_{4},x_{5},x_{6},x_{0}^{2}+x_{1}^{2}+3x_{2}^{2})$ and $L_{56}=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C$ is a smooth conic (degree $2$) and $L$ is a line (degree $1$). We have $C\cap L=\varnothing$.
@@ 5.20 Case $k=20$
Let $X_{20}:=V(\phi^{\mathrm{nf}}_{20})\subset\mathbb{P}^{6}$. Then the singular locus consists of two disjoint components:
@@
where $C=V(x_{4},x_{5},x_{6},4x_{1}^{2}-2x_{2}^{2}+x_{3}^{2},4x_{0}^{2}+6x_{2}^{2}+3x_{3}^{2})$ and $L=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C$ is a curve of degree $4$ (a complete intersection of two quadrics, which is an elliptic curve) and $L$ is a line (degree $1$). We have $C\cap L=\varnothing$.
@@ 1 $C\cup\{P\}$ ($C$ smooth conic) conic (deg 2) one point $P$ of type $\mathsf{E}_{19}$
2 $L_{01}\cup C$ ($L_{01}\cong\mathbb{P}^{1}$, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 4)
3 $L_{01}\cup C$ ($L_{01}\cong\mathbb{P}^{1}$, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 4)
@@ 9 $L_{01}\cup\{P\}$ ($L_{01}$ line) line (deg 1) one point $P$ of type $\mathcal{O}_{6}$
10 $L_{01}\cup C$ ($L_{01}$ line, $C$ conic) line ++ conic (deg 2)
@@ 11 $\{P_{1}\}\cup\{P_{2}\}\cup\{P_{3}\}$ three points three points $P_{i}$ of type $\mathcal{C}_{cub}$
12 $L_{56}\cup C$ ($L_{56}$ line, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 4)
13 $C_{1}\cup C_{2}$ (disjoint) two pairs of intersecting lines
14 $L_{01}\cup C$ ($L_{01}$ line, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 4)
15 $C_{1}\cup C_{2}\cup L$ (two CI(2,2)(2,2)s + line) two quartics ++ line
@@ 16 $\{P\}\cup\{Q\}$ two points $P,Q$ of type $\mathsf{E}_{20}$ (same modulus)
17 $C\cup L$ ($C$ smooth conic, $L$ line) conic (deg 2) ++ line
18 $L_{1}\cup L_{2}$ (two disjoint lines) two disjoint lines
@@ 19 $C\cup\{P\}$ ($C$ smooth conic) conic (deg 2) one point $P$ of type $\mathsf{E}_{18}$
20 $C\cup L$ ($C$ elliptic curve, $L$ line) elliptic curve (deg 4) ++ line
@@ 21 $\{P_{1}\}\cup\{P_{2}\}\cup\{P_{3}\}$ three points three points $P_{i}$ of type $\mathcal{C}_{cub}$
$\mathsf{E}_{20}$ $\mathrm{QH}(1,2,2;6)_{20}(t)$ $Y^{3}+Z^{3}+X^{6}+t_{1}X^{2}Y^{2}+\cdots+t_{5}X^{4}Z=0$
5.22 Minimal exponents and connections to Park’s stability criterion
In Definitions 5.1–5.4 and 5.6–5.7 we introduced six quasi-homogeneous isolated hypersurface singularities that occur as the isolated points on the GIT boundary. The next theorem records a uniform numerical feature of these singularities, and explains why they sit exactly at the critical value in Park’s stability criterion.
@@ 5.22 Minimal exponents and connections to Park’s stability criterion
where $c=\mathrm{corank}_{x}(X)$ and $g$ is one of the weighted-homogeneous germs from Definitions 5.1–5.4 and 5.6–5.7.
By Thom–Sebastiani for minimal exponents [Par25, Theorem 2.3],
@@
Significance (relation to Park). Since $\frac{7}{3}>2$, these isolated singularities are $1$-rational (hence in particular rational and Du Bois) by [Par25, Theorem 2.2]. More importantly, for cubic fivefolds $(n,d)=(6,3)$ we have $\frac{n+1}{d}=\frac{7}{3}$, which is exactly the threshold in Park’s GIT criterion: a degree $d$ hypersurface $X\subset\mathbb{P}^{n}$ is GIT stable (resp. semistable) if $\widetilde{\alpha}(X)>\frac{n+1}{d}$ (resp. $\geq$) [Par25, Theorem A]. Thus, any cubic fivefold containing one of these singular points satisfies $\widetilde{\alpha}(X)\leq\frac{7}{3}$, so it cannot lie in the region $\widetilde{\alpha}(X)>\frac{7}{3}$ guaranteed to be GIT stable by Park. This gives a conceptual explanation for why these singularities occur precisely on the GIT boundary.
Definition 5.30.
We call the above six singularity types extremal cubic fivefold singularities, since they realize the critical value
$\frac{n+1}{d}=\frac{7}{3}\qquad(n=6,\ d=3)$
in Park’s GIT criterion. In particular, each of these types has local minimal exponent $\frac{7}{3}$. Motivated by this, we refer to isolated cubic fivefold singularities with local minimal exponent $\frac{7}{3}$ as $\frac{7}{3}$-singularities; the six types above occur on the GIT boundary.
Remark 5.31 (Unification via minimal exponent).
@@
Park observes that, for cubic hypersurfaces $X\subset\mathbb{P}^{n}$ with $n\leq 5$, the classical explicit GIT analyses can be reformulated in terms of the same numerical threshold $\frac{n+1}{3}$ for the (global) minimal exponent; see [Par25, Remark 7.6]. Theorem 5.29 shows that for cubic fivefolds $(n=6)$ the isolated boundary singularities also satisfy equality $\widetilde{\alpha}_{x}(X)=\frac{7}{3}=\frac{n+1}{3}$, fitting the same pattern.
@@ 6 Adjacency relations among strictly semistable components
In this section, we record the adjacency relations among the closed strata $\{\Phi_{k}\}_{k=1}^{21}\subset\mathbb{P}(W)^{\mathrm{ss}}$, where $W=\mathrm{Sym}^{3}\mathbb{C}^{7}$. We adopt the following convention: two closed strata $\Phi_{i}$ and $\Phi_{j}$ are adjacent if there exists a codimension-one wall in the Hilbert–Mumford weight space such that the maximally destabilizing one-parameter subgroups (1-PS’s) for general points of $\Phi_{i}$ and $\Phi_{j}$ coincide on that wall, and the corresponding degenerations specialize to the same unstable limit in $\mathbb{P}(W)$ (see Definition 6.2).
This notion of adjacency captures wall-crossing in Kirwan’s stratification and, from the combinatorial viewpoint, corresponds to codimension-one faces of the cones $I(r)_{\geq 0}$ determined by 1-PS weight vectors $r$.
Our strategy is as follows.

### v7

@@ 5.14 Case $k=14$
Let $X_{14}:=V(\phi^{\mathrm{nf}}_{14}(\tau))\subset\mathbb{P}^{6}$. Then the singular locus has two irreducible components:
@@
where $L_{01}=V(x_{2},x_{3},x_{4},x_{5},x_{6})$ and
$C=V(x_{0},x_{1},x_{2},(\tau-1)x_{4}^{2}-2x_{5}^{2}+x_{6}^{2},(\tau-1)x_{3}^{2}+(\tau+1)x_{5}^{2}+(\tau-2)x_{6}^{2}).$
We have $L_{01}\cap C=\varnothing$.
@@ 5.15 Case $k=15$
Let $X_{15}:=V(\phi^{\mathrm{nf}}_{15}(\tau,\sigma))\subset\mathbb{P}^{6}$. Then the singular locus consists of three irreducible components:
@@
where $C_{1}=V(x_{1},x_{3},x_{6},\tau x_{4}^{2}+x_{0}x_{5},\sigma x_{2}^{2}+x_{0}x_{4})$, $C_{2}=V(x_{2},x_{4},x_{5},x_{3}^{2}+x_{0}x_{6},x_{1}^{2}+x_{0}x_{3})$, and $L=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C_{1}$ and $C_{2}$ are curves of degree $4$ (complete intersections of two quadrics in $\mathbb{P}^{3}$), and $L$ is a line.
@@ 5.17 Case $k=17$
Let $X_{17}:=V(\phi^{\mathrm{nf}}_{17}(\tau,\sigma))\subset\mathbb{P}^{6}$. Then the singular locus consists of two disjoint components:
@@
where $C=V(x_{3},x_{4},x_{5},x_{6},x_{0}^{2}+x_{1}^{2}+\sigma x_{2}^{2})$ and $L_{56}=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C$ is a smooth conic (degree $2$) and $L$ is a line (degree $1$). We have $C\cap L_{56}=\varnothing$.
@@ 5.20 Case $k=20$
Let $X_{20}:=V(\phi^{\mathrm{nf}}_{20}(\tau))\subset\mathbb{P}^{6}$. Then the singular locus consists of two disjoint components:
@@
where $C=V(x_{4},x_{5},x_{6},(\tau-1)x_{1}^{2}-2x_{2}^{2}+x_{3}^{2},(\tau-1)x_{0}^{2}+(\tau+1)x_{2}^{2}+(\tau-2)x_{3}^{2})$ and $L=V(x_{0},x_{1},x_{2},x_{3},x_{4})$. Here, $C$ is a curve of degree $4$ (a complete intersection of two quadrics, which is an elliptic curve) and $L$ is a line (degree $1$). We have $C\cap L=\varnothing$.
@@ 1 $C\cup\{P\}$ ($C$ smooth conic) conic (deg 2) one point $P$ of type $\mathsf{E}_{19}$
2 $L_{01}\cup C$ ($L_{01}\cong\mathbb{P}^{1}$, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 1+4)
3 $L_{01}\cup C$ ($L_{01}\cong\mathbb{P}^{1}$, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 1+4)
@@ 9 $L_{01}\cup\{P\}$ ($L_{01}$ line) line (deg 1) one point $P$ of type $\mathcal{O}_{6}$
10 $L_{01}\cup C$ ($L_{01}$ line, $C$ conic) line ++ conic (deg 1+2)
@@ 11 $\{P_{1}\}\cup\{P_{2}\}\cup\{P_{3}\}$ three points three points $P_{i}$ of type $\mathcal{C}_{cub}$
12 $L_{56}\cup C$ ($L_{56}$ line, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 1+4)
13 $C_{1}\cup C_{2}$ (disjoint) two pairs of intersecting lines (deg 1+1+1+1)
14 $L_{01}\cup C$ ($L_{01}$ line, $C$ a CI(2,2)(2,2)) line ++ quartic curve (deg 1+4)
15 $C_{1}\cup C_{2}\cup L$ (two CI(2,2)(2,2)s + line) two quartics ++ line (deg 4+4+1)
@@ 16 $\{P\}\cup\{Q\}$ two points $P,Q$ of type $\mathsf{E}_{20}$ (same modulus)
17 $C\cup L$ ($C$ smooth conic, $L$ line) conic ++ line (deg 2+1)
18 $L_{1}\cup L_{2}$ (two disjoint lines) two disjoint lines (deg 1+1)
@@ 19 $C\cup\{P\}$ ($C$ smooth conic) conic (deg 2) one point $P$ of type $\mathsf{E}_{18}$
20 $C\cup L$ ($C$ elliptic curve, $L$ line) elliptic curve ++ line (deg 4+1)
@@ 21 $\{P_{1}\}\cup\{P_{2}\}\cup\{P_{3}\}$ three points three points $P_{i}$ of type $\mathcal{C}_{cub}$
$\mathsf{E}_{20}$ $\mathrm{QH}(1,2,2;6)_{20}(t)$ $\scriptstyle{Y^{3}+Z^{3}+X^{6}+t_{1}X^{2}Y^{2}+t_{2}X^{2}YZ+t_{3}Y^{2}Z+t_{4}X^{4}Y+t_{5}X^{4}Z=0}$
6 The minimal exponents
6.1 Minimal exponents of isolated singularities
In Definitions 5.1–5.6 we introduced six quasi-homogeneous isolated hypersurface singularities that occur as the isolated points on the GIT boundary. The next theorem records a uniform numerical feature of these singularities, and explains why they sit exactly at the critical value in Park’s stability criterion.
@@ 6.1 Minimal exponents of isolated singularities
where $c=\mathrm{corank}_{x}(X)$ and $g$ is one of the weighted-homogeneous germs from Definitions 5.1–5.6
By the Thom-Sebastiani type theorem for minimal exponents (see, for example, [Par25, Theorem 2.3]),
@@
For cubic fivefolds $(n,d)=(6,3)$, Park’s GIT criterion gives the sharp threshold $\frac{n+1}{d}=\frac{7}{3}$: a degree-$d$ hypersurface $X\subset\mathbb{P}^{n}$ is GIT stable (resp. semistable) if $\widetilde{\alpha}(X)>\frac{n+1}{d}$ (resp. $\geq\frac{n+1}{d}$) [Par25, Theorem A]. Hence, if $X$ has a point with local minimal exponent $7/3$, then $\widetilde{\alpha}(X)\leq 7/3$, so $X$ cannot lie in the strictly stable locus. This explains why the isolated singularities in Theorem 6.1 occur precisely on the GIT boundary.
@@
Moreover, Park observed that the classical explicit GIT analyses for cubic hypersurfaces with $n\leq 5$ can be reformulated uniformly in terms of the same threshold $\frac{n+1}{3}$ for the global minimal exponent [Par25, Remark 7.6]; Theorem 6.1 shows that, for cubic fivefolds ($n=6$), the equality
$\widetilde{\alpha}_{x}(X)=\frac{7}{3}=\frac{n+1}{3}$
is attained along the strictly semistable boundary.
Definition 6.3 (Extremal and $7/3$-singularities).
We call the six isolated analytic types appearing in Theorem 6.1 extremal cubic fivefold singularities. More generally, an isolated cubic fivefold singularity with local minimal exponent $7/3$ is called a $7/3$-singularity.
6.2 The global minimal exponents of general members of the boundary families
For families whose singular locus $\mathrm{Sing}(X_{k})$ has positive-dimensional components (curves or lines), we must compute the local minimal exponent both at a general point and at the special points along these strata in order to determine the global minimal exponent
$\widetilde{\alpha}(X_{k})\;=\;\min_{x\in\mathrm{Sing}(X_{k})}\widetilde{\alpha}_{x}(X_{k}).$
Our computations for all $21$ closed-orbit normal forms reveal a striking uniformity that matches Park’s threshold:
Theorem 6.4.
For each $k\in\{1,\dots,21\}$, the global minimal exponent of a general strictly semistable closed-orbit representative $X_{k}=V(\phi^{\mathrm{nf}}_{k})\subset\mathbb{P}^{6}$ in the component $\Phi_{k}$ is
$\widetilde{\alpha}(X_{k})\;=\;\frac{7}{3}.$
Proof.
We prove the theorem by a uniform case analysis organized by the structure of the one-dimensional components of $\mathrm{Sing}(X_{k})$. More precisely, for each $k$ with $\dim\mathrm{Sing}(X_{k})=1$, we classify every irreducible curve component $\Gamma\subset\mathrm{Sing}(X_{k})$ into one of finitely many pattern types according to the behavior of the local minimal exponent $p\mapsto\widetilde{\alpha}_{p}(X_{k})$ along $\Gamma$ (see Table 5).
In each pattern type, the minimum of $\widetilde{\alpha}_{p}(X_{k})$ along $\Gamma$ equals $7/3$. For types A1 and A2, this minimum is attained at an explicitly described finite subset of special points, whereas for types B and C, the minimal exponent is identically $7/3$ everywhere along $\Gamma$. Consequently, $\min_{p\in\Gamma}\widetilde{\alpha}_{p}(X_{k})=\frac{7}{3}$ for every curve component $\Gamma\subset\mathrm{Sing}(X_{k})$.
If $\mathrm{Sing}(X_{k})$ has only isolated singular points (namely $k\in\{5,11,16,21\}$), then $\widetilde{\alpha}(X_{k})=7/3$ follows directly from the computation of the six isolated extremal singularity types recorded in Table 4. Taking the minimum over all singular points, we obtain $\widetilde{\alpha}(X_{k})=7/3$ for every $k=1,\dots,21$. ∎
component type Type A1 Type A2 Type B Type C
line
$k=2$: $L_{01}$
$k=3$: $L_{01}$
$k=4$: $L_{01},\,L_{56}$
$k=6$: $L_{56}$
$k=10$: $L_{01}$
$k=12$: $L_{56}$
$k=17$: $L_{56}$
$k=18$: $L_{01},\,L_{56}$
$k=9$: $L_{01}$
$k=14$: $L_{01}$
$k=15$: $L$
$k=20$: $L$
smooth conic
$k=1$: $C$
$k=7$: $C$
$k=10$: $C$
$k=17$: $C$
$k=8$: $C$
$k=19$: $C$
twisted quartic
$k=6$: $C$
$k=12$: $C$
elliptic quartic
$k=14$: $C$
$k=20$: $C$
reducible conic $k=13$: $C_{1},\,C_{2}$
CI(2,2)
$k=2$: $C$
$k=3$: $C$
$k=15$: $C_{1},\,C_{2}$
Table 5: Index of one-dimensional singular-locus components by minimal-exponent pattern type.
Pattern types for one-dimensional singular components
Let $X\subset\mathbb{P}^{6}$ be one of the closed-orbit normal forms $X_{k}=V(\phi^{\mathrm{nf}}_{k})$, and let $\Gamma\subset\mathrm{Sing}(X)$ be an irreducible curve component. We classify $\Gamma$ according to the behavior of the function
$\Gamma\ni p\longmapsto\widetilde{\alpha}_{p}(X)$
along $\Gamma$. In all cases appearing in this paper, the only values that occur on $\Gamma$ are $5/2$ and $7/3$. We say that $\Gamma$ is of one of the following pattern types.
Type A1 (one special point). $\Gamma$ is of Type A1 if there exists a unique point $p_{0}\in\Gamma$ such that
$\widetilde{\alpha}_{p_{0}}(X)=\frac{7}{3},\qquad\widetilde{\alpha}_{p}(X)=\frac{5}{2}\ \text{ for all }p\in\Gamma\setminus\{p_{0}\}.$
Equivalently, $\widetilde{\alpha}_{p}(X)=5/2$ at a general point of $\Gamma$, and the minimum along $\Gamma$ is attained at a single explicitly described “special” point $p_{0}$.
Type A2 (two special points). $\Gamma$ is of Type A2 if there exist exactly two points $p_{1},p_{2}\in\Gamma$ such that
$\widetilde{\alpha}_{p_{1}}(X)=\widetilde{\alpha}_{p_{2}}(X)=\frac{7}{3},\qquad\widetilde{\alpha}_{p}(X)=\frac{5}{2}\ \text{ for all }p\in\Gamma\setminus\{p_{1},p_{2}\}.$
Equivalently, $\widetilde{\alpha}_{p}(X)=5/2$ at a general point of $\Gamma$, and the minimum along $\Gamma$ is attained at two explicitly described special points.
Type B (uniform $7/3$ without rank drop). $\Gamma$ is of Type B if
$\widetilde{\alpha}_{p}(X)=\frac{7}{3}\quad\text{for all }p\in\Gamma,$
and moreover the transverse quadratic part to $\Gamma$ has constant rank along $\Gamma$ (i.e. there are no rank-drop points of the Hessian along $\Gamma$). In particular, the minimum of $\widetilde{\alpha}_{p}(X)$ on $\Gamma$ equals $7/3$ and is attained everywhere on $\Gamma$.
Type C (uniform $7/3$ with rank drop). $\Gamma$ is of Type C if
$\widetilde{\alpha}_{p}(X)=\frac{7}{3}\quad\text{for all }p\in\Gamma,$
but the transverse quadratic part degenerates at finitely many points of $\Gamma$ (equivalently, the Hessian rank drops along $\Gamma$ at a finite set). Thus Type C shares the same uniform value $7/3$ as Type B, but differs in that certain special points on $\Gamma$ exhibit additional quadratic degenerations, which nevertheless do not change the minimal exponent.
In all cases in this paper, every curve component $\Gamma\subset\mathrm{Sing}(X_{k})$ is of one of the types A1, A2, B, or C; see Table 5.
@@ 7 Adjacency relations among strictly semistable components
In this section we record the codimension-one wall adjacency among the closed strata $\{\Phi_{k}\}_{k=1}^{21}\subset\mathbb{P}(W)^{\mathrm{ss}}$, where $W=\mathrm{Sym}^{3}\mathbb{C}^{7}$. Throughout Section 7, we use the term “adjacent” only in the sense of wall adjacency (i.e. codimension-one incidence in the Hilbert–Mumford weight fan).
Sign convention and why we match $\pm r_{j}$. A wall is determined by a hyperplane, hence it is insensitive to the orientation of a normal vector. Consequently, when we compare slices of the form
$I(r)_{\geq 0}\cap I(r_{k})_{=0}\subset I(r_{k})_{=0},$
the same geometric wall may be represented either by $r$ or by r-r. For this reason, in the computational matching against the reduced list $r_{1},\dots,r_{21}$ (Section 2), we always test both signs $\pm r_{j}$. This eliminates the spurious “one-way” phenomena that occur if one only matches the fixed sign representatives $r_{j}$.
For background on wall-crossing in GIT and Kirwan’s stratification, see [Kir84, DH98, Tha96].
Definition 7.1 (Adjacency via codimension-one walls).
Let $\Phi_{i},\Phi_{j}$ be closed strata in the strictly semistable locus. We say that $\Phi_{i}$ and $\Phi_{j}$ are adjacent if the associated maximal cones $I(r_{i})_{\geq 0}$ and $I(r_{j})_{\geq 0}$ meet along a codimension-one face (a wall) in the Hilbert–Mumford weight fan. Equivalently, there exists a codimension-one wall on which the Hilbert–Mumford numerical functions for the corresponding maximally destabilizing $1$-PS’s tie, and the wall separates the chambers corresponding to $r_{i}$ and $r_{j}$.
7.1 Maximal slices and the computational criterion
Fix an index $k$. Inside the hyperplane $I(r_{k})_{=0}\subset I$ we consider subsets of the form
$I(r)_{\geq 0}\cap I(r_{k})_{=0}\qquad(r\in\mathbb{Z}^{7}_{(0)},\ r\not\parallel r_{k}).$
Such a subset corresponds to a codimension-one face of the cone $I(r_{k})_{\geq 0}$ precisely when it is inclusion-maximal among all subsets of this form. Following the terminology used in the accompanying Sage computations, we call these inclusion-maximal subsets the maximal slices in $I(r_{k})_{=0}$.
Directed slice-matching relation. For bookkeeping, we introduce a directed relation on $\{1,\dots,21\}$: we write $k\rightsquigarrow j$ if some maximal slice in $I(r_{k})_{=0}$ equals
$I(\pm r_{j})_{\geq 0}\cap I(r_{k})_{=0}$
for at least one choice of sign. (Here $\pm r_{j}$ indicates that we match against both orientations of $r_{j}$, as explained above.)
By Definition 7.1, wall adjacency is symmetric. In our combinatorial model this means:
$\Phi_{i}\text{ and }\Phi_{j}\text{ are adjacent}\quad\Longleftrightarrow\quad(i\rightsquigarrow j)\ \text{and}\ (j\rightsquigarrow i).$
In other words, the adjacency graph is obtained from the directed slice-matching digraph by retaining only the mutual pairs.
Implementation. We compute maximal slices by the same linear-algebraic method as in Section 2, adapted to the codimension-one setting: we enumerate $4$-tuples of points in $I(r_{k})_{=0}$, extract candidate normals from the right kernel (modulo $r_{k}$), build the corresponding slices, and keep only inclusion-maximal ones. We then match each maximal slice against the list $\{\pm r_{1},\dots,\pm r_{21}\}$. (See the accompanying Sage scripts and logs for the full output.)
7.2 Codimension-one adjacency graph
The main output of the codimension-one computation is the following.
Theorem 7.2 (Codimension-one wall adjacency).
Among the closed strata $\Phi_{k}$ in the strictly semistable locus, the wall-adjacency graph (codimension-one adjacency in the sense of Definition 7.1) has $21$ vertices and $56$ edges. Equivalently, for each $k$ the set
$N(k)\ :=\ \{\,j\neq k\mid\Phi_{k}\text{ is adjacent to }\Phi_{j}\,\}$
is given by Table 6.
Table 6: Codimension-one wall adjacency: for each $k$, the neighbor set $N(k)$ of indices $j$ such that $\Phi_{k}$ and $\Phi_{j}$ are adjacent (Definition 7.1).
$k$ $N(k)$ $k$ $N(k)$
11 $3,5,7,10$ $12$ $3,4,5,6,7,15,17$
22 $3,4,5,9$ $13$ $4,8,18,19$
33 $1,2,4,5,9,10,12$ $14$ $9,10,11,18$
44 $2,3,5,6,12,13,18$ $15$ $6,8,12,17,18,20$
55 $1,2,3,4,6,7,12$ $16$ $8,18,19$
66 $4,5,12,15$ $17$ $7,12,15,18,19,20,21$
77 $1,5,12,17$ $18$ $4,9,10,13,14,15,16,17,20$
88 $10,11,13,15,16,19$ 1919 $8,9,13,16,17,21$
99 $2,3,10,14,18,19$ 2020 $15,17,18,21$
$10$ $1,3,8,9,11,14,18$ 2121 $17,19,20$
$11$ $8,10,14$
Figure 2: Codimension-one slice-matching graph for the strata $\Phi_{1},\dots,\Phi_{21}$. Vertices are indexed by $k$ and represent $\Phi_{k}$. A dotted arrow $i\to j$ indicates a one-sided slice match: there exists an inclusion-maximal slice $S\subset I(r_{i})_{=0}$ of the form $S=I(\pm r_{j})_{\geq 0}\cap I(r_{i})_{=0}$, while no such match occurs in the reverse direction. A solid (undirected) edge between $i$ and $j$ indicates that slice matches occur in both directions ($i\to j$ and $j\to i$); equivalently, $\Phi_{i}$ and $\Phi_{j}$ are adjacent in the sense of Definition 7.1.
Remark 7.3 (One-sided slice matches and higher codimension).
In the directed slice-matching digraph one also encounters one-sided arrows $k\rightsquigarrow j$ without the reverse arrow $j\rightsquigarrow k$. These do not represent codimension-one wall adjacency in the sense of Definition 7.1; rather, they indicate that a facet of $I(r_{k})_{\geq 0}$ is detected by $\pm r_{j}$ when restricted to $I(r_{k})_{=0}$, while from the $j$-side the same incidence is not codimension-one. Empirically, such one-sided phenomena are accounted for by higher-codimension incidence among faces (codimension $\geq 2$), as can be explored by extending the same slice computation to intersections of multiple hyperplanes $I(r_{k_{1}})_{=0}\cap\cdots\cap I(r_{k_{c}})_{=0}$.
8 Non-inclusions and Filters
In this section, we prove that there are no inclusion relations among the $21$ boundary families of strictly semistable cubic fivefolds constructed in the previous sections. Concretely, for any distinct indices $k,\ell\in\{1,\dots,21\}$ we show
$\Phi_{k}\not\subseteq\Phi_{\ell}.$
Equivalently, we must rule out all $21\cdot 20=420$ ordered pairs $(k,\ell)$ with $k\neq\ell$.
Rather than checking these pairs directly, we use a sieve (or “filter”) method: we impose a sequence of necessary conditions for an inclusion $\Phi_{k}\subseteq\Phi_{\ell}$. Each condition is monotone under specialization (typically by upper semicontinuity), so any pair violating the condition is discarded. Applying these filters successively shrinks the candidate list until it becomes empty.
The filters are applied in the following order.